### v2

# Dichotomies uniform on subspaces and formulas for dichotomy spectra

Abstract
We introduce a notion of dichotomy which generalizes the classical concept of exponential dichotomy and the recent notion of Bohl dichotomy. A key attribute is the discussion of the sets of subspaces of the state space on which the dichotomy estimates are uniform. Two main results are a dichotomy spectral theorem based on our notion of dichotomy which is uniform on subspaces and a formula for the dichotomy spectral intervals which is new for the Bohl dichotomy spectrum as well as for the classical exponential dichotomy spectrum.
Keywords: hyperbolicity, exponential dichotomy, Bohl dichotomy, spectral theorem
1 Introduction
In this paper we develop formulas for the dichotomy spectrum of linear time-varying difference equations on $\mathbb{R}^{d}$, $d\in\mathbb{N}_{>0}$
$x(n+1)=A(n)x(n),\qquad n\in\mathbb{T},$ (1)
with one-sided time $\mathbb{T}=\mathbb{N}$ or two-sided time $\mathbb{T}=\mathbb{Z}$ and invertible coefficients $A\colon\mathbb{T}\to\mathrm{GL}(\mathbb{R}^{d})$ such that $A$ and $A^{-1}\colon\mathbb{T}\to\mathrm{GL}(\mathbb{R}^{d})$, $n\mapsto A(n)^{-1}$ are bounded.
There is a rich body of literature on hyperbolicity concepts for systems (1), most prominently exponential dichotomy, but also more recently Bohl dichotomy, see Definition 1 and [20] (for the case $\mathbb{T}=\mathbb{N}$) and Remark 2 for a short history and references. In this introduction we observe that a main difference between these known dichotomy notions is the set of subspaces on which the dichotomy estimates for the solutions of (1) are uniform. This will lead us in Definition 4 to a new notion of dichotomy which is uniform on subspaces and which generalizes the concepts of exponential and Bohl dichotomy.
Section 2 contains results on properties of dichotomous systems (1) as a preparation for the main results in Section 4. In Section 3 we recall in Definition 14 Bohl exponents from [20] and investigate their relation to our general notion of dichotomy. Section 4 contains the two main results, a dichotomy spectral theorem based on our general notion of dichotomy which is uniform on subspaces (Theorem 29) and a formula for the dichotomy spectral intervals (Theorem 30).
For an initial value $x_{0}\in\mathbb{R}^{d}$ the solution $x(\cdot,x_{0}):\mathbb{T}\to\mathbb{R}^{d}$ of (1) starting at 0 in $x_{0}$ satisfies $x(0,x_{0})=x_{0}$,
$x(n,x_{0})=A(n-1)\cdots A(0)x_{0},\qquad n>0,$
and if $\mathbb{T}=\mathbb{Z}$
$x(n,x_{0})=A(n)^{-1}\cdots A(-1)^{-1}x_{0},\qquad n<0.$
Dichotomies are solution estimates on subspaces of the state space $\mathbb{R}^{d}$ which form a splitting. Denote for a subspace $L\subseteq\mathbb{R}^{d}$ the set of subspaces of $L$ by
$\mathcal{G}(L)$ $\coloneqq\{U\subseteq L:\text{$U$ subspace}\}.$
For $L_{1},L_{2}\in\mathcal{G}(\mathbb{R}^{d})$ we say that $(L_{1},L_{2})$ is a splitting of $\mathbb{R}^{d}$ if $L_{1}\oplus L_{2}=\mathbb{R}^{d}$.
Let $(L_{1},L_{2})$ be a splitting of $\mathbb{R}^{d}$. We denote by $\pi_{L_{1}}$ the projection onto $L_{1}$ along $L_{2}$. Then $\pi_{L_{2}}\coloneqq I-\pi_{L_{1}}$ is the projection onto $L_{2}$ along $L_{1}$, where $I$ denotes the identity mapping on $\mathbb{R}^{d}$. For $V\in\mathcal{G}(\mathbb{R}^{d})$ and $i\in\{1,2\}$
$\pi_{L_{i}}[V]\coloneqq\{\pi_{L_{i}}v:v\in V\}$
is the projection of $V$ onto $L_{i}$. For the following definition see e.g. [20] (or [19]; note that system (1) has bounded growth) and the references therein.
Definition 1 (Exponential and Bohl dichotomy).
Let $(L_{1},L_{2})$ be a splitting of $\mathbb{R}^{d}$. System (1) has a Bohl dichotomy on $(L_{1},L_{2})$ if there exists an $\alpha>0$ and functions $C_{1},C_{2}:\mathbb{R}^{d}\to(0,\infty)$ such that
$\|x(n,x_{0})\|$ $\leq C_{1}(x_{0})e^{-\alpha(n-m)}\|x(m,x_{0})\|,$ $\quad m,n\in\mathbb{T},n\geq m,x_{0}\in L_{1},$
$\|x(n,x_{0})\|$ $\geq C_{2}(x_{0})e^{\alpha(n-m)}\|x(m,x_{0})\|,$ $\quad m,n\in\mathbb{T},n\geq m,x_{0}\in L_{2}.$
In case $C_{1}$ and $C_{2}$ can be chosen independent of $x_{0}$, system (1) has an exponential dichotomy.
Remark 2 (History of dichotomy notion).
(a) Foundations. Massera and Schaeffer [31, 32] coined the name exponential dichotomy and formulated the canonical definition of exponentially dichotomous linear differential equations. The foundational formulation of the theory of exponentially dichotomous systems was completed through the monographs of Massera and Schaeffer [30] and Daleckii and Krein [22], which summarized earlier results on exponentially dichotomous systems and outlined paths for the development of this theory by presenting numerous open questions.
For difference equations, this concept first appeared in the work of Coffman and Schaeffer [18], where the notion of uniform exponential dichotomy for infinite-dimensional, one-sided discrete-time systems was introduced, highlighting differences between continuous and discrete-time systems due to the potential non-invertibility of the transition operator. The first definitions of dichotomy for two-sided sequences were provided in [23] and [26, Definition 7.6.4]. Subsequently, many authors have studied the concept of uniform exponential dichotomies for discrete systems with both invertible and non-invertible coefficients [14, 15, 1, 25].
The origins of the concept go back to Perron’s seminal work [35], one of the earliest studies of this property. For comprehensive treatments of exponential dichotomy and its applications, see [19, 29, 34].
(b) Concepts of uniformity. The term uniform usually refers to the fact that the dichotomy estimates are uniform with respect to both time and the initial condition. While the role of uniformity with respect to time has been extensively analyzed in the literature (see e.g. [10, 11, 12, 43] and references therein), research of the significance of uniformity with respect to the initial condition has only recently been initiated (see e.g. [13, 24, 9, 20, 21]). These studies examine a type of dichotomy, referred to as Bohl dichotomy in [20], where the convergence with respect to time is uniform, but uniformity with respect to the initial condition is restricted to each one-dimensional subspace.
(c) Weak dichotomies. The version of exponential dichotomy where the constants $C_{1}$, $C_{2}$ in Definition 1 may depend on the initial condition was first introduced for continuous-time systems in [13], without a specific name. It was later explicitly termed weak exponential dichotomy in [6]. For discrete-time systems on the half-line, the same notion appeared in [9] and has since been systematically studied under the name Bohl dichotomy in [20, 21].
In the following remark we emphasize the subspaces on which the dichotomy estimates in Definition 1 hold uniformly. Denote for $k\in\mathbb{N}$ and a subspace $L\subseteq\mathbb{R}^{d}$ the set of $k$-dimensional subspaces of $L$ by
$\mathcal{G}_{k}(L)$ $\coloneqq\{U\in\mathcal{G}(L):\dim U=k\}.$
Remark 3 (Uniformity subspaces of Bohl and exponential dichotomy).
If (1) has a Bohl dichotomy on $(L_{1},L_{2})$ then, using the homogeneity of the solution $x(\cdot,cx_{0})=cx(\cdot,x_{0})$, $c\in\mathbb{R}$, the dichotomy estimates hold uniformly on all one-dimensional subspaces of $L_{1}$ and $L_{2}$, respectively. More precisely, for all $U\in\mathcal{G}_{1}(L_{1})$, there exists $C(U)>0$ with
$\|x(n,x_{0})\|\leq C(U)\mathrm{e}^{-\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{T},\,n\geq m,\,x_{0}\in U,$
and for all $U\in\mathcal{G}_{1}(L_{2})$, there exists $C(U)>0$ with
$\|x(n,x_{0})\|\geq C(U)\mathrm{e}^{\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{T},\,n\geq m,\,x_{0}\in U.$
If (1) has an exponential dichotomy on $(L_{1},L_{2})$ then the dichotomy estimates hold uniformly on all subspaces of $L_{1}$ and $L_{2}$, respectively.
We introduce a new dichotomy notion based on dichotomy estimates which hold uniformly on a set of subspaces of the splitting which covers the splitting.
Definition 4 (Dichotomy uniform on subspaces of a splitting).
Let $(L_{1},L_{2})$ be a splitting of $\mathbb{R}^{d}$ and $\mathcal{U}_{1}\subseteq\mathcal{G}(L_{1})$, $\mathcal{U}_{2}\subseteq\mathcal{G}(L_{2})$ be covers of $L_{1}$ and $L_{2}$, respectively, i.e.
$\bigcup\mathcal{U}_{1}=L_{1}\qquad\text{and}\qquad\bigcup\mathcal{U}_{2}=L_{2}.$ (2)
We say that system (1) has a dichotomy on $(L_{1},L_{2})$ uniformly on the subspaces in $(\mathcal{U}_{1},\mathcal{U}_{2})$ if there exists an $\alpha>0$ such that for each $U_{1}\in\mathcal{U}_{1}$, there exists $C(U_{1})>0$ with
$\|x(n,x_{0})\|\leq C(U_{1})\mathrm{e}^{-\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{T},\,n\geq m,\,x_{0}\in U_{1},$ (3)
and for each $U_{2}\in\mathcal{U}_{2}$, there exists $C(U_{2})>0$ with
$\|x(n,x_{0})\|\geq C(U_{2})\mathrm{e}^{\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{T},\,n\geq m,\,x_{0}\in U_{2}.$ (4)
We say that system (1) has a dichotomy on $(L_{1},L_{2})$ if there exist covers $\mathcal{U}_{1}\subseteq\mathcal{G}(L_{1})$ of $L_{1}$ and $\mathcal{U}_{2}\subseteq\mathcal{G}(L_{2})$ of $L_{2}$ such that system (1) has a dichotomy on $(L_{1},L_{2})$ uniformly on the subspaces in $(\mathcal{U}_{1},\mathcal{U}_{2})$.
Remark 5 (Examples of dichotomies uniform on subspaces of a splitting).
By Definition 1 Bohl and exponential dichotomy are examples of dichotomies uniform on subspaces. If system (1) has a dichotomy uniform on subspaces $(\mathcal{U}_{1},\mathcal{U}_{2})$ of a splitting $(L_{1},L_{2})$ then it has an
(a) exponential dichotomy if $\mathcal{U}_{1}=\{L_{1}\}$ and $\mathcal{U}_{2}=\{L_{2}\}$ and a
(b) Bohl dichotomy if $\mathcal{U}_{1}=\mathcal{G}_{1}(L_{1})$ and $\mathcal{U}_{2}=\mathcal{G}_{1}(L_{2})$.
The important and obvious observation that a dichotomy estimate which holds uniformly on a subspace $U$ also holds uniformly on all subspaces of $U$ is formulated in the following theorem.
Theorem 6 (Refining uniformity subspaces of dichotomy).
If system (1) has a dichotomy on $(L_{1},L_{2})$ uniformly on the subspaces in $(\mathcal{U}_{1},\mathcal{U}_{2})$, then it also has a dichotomy on $(L_{1},L_{2})$ uniformly on the subspaces in the refinement $(\mathcal{V}_{1},\mathcal{V}_{2})$ with
$\mathcal{V}_{1}\coloneqq\bigcup_{U\in\mathcal{U}_{1}}\mathcal{G}(U)\qquad\text{and}\qquad\mathcal{V}_{2}\coloneqq\bigcup_{U\in\mathcal{U}_{2}}\mathcal{G}(U).$
Proof.
Let $i\in\{1,2\}$. The set $\mathcal{V}_{i}$ consists of all subspaces of the subspaces in $\mathcal{U}_{i}$ and therefore it also covers $L_{i}$. Let $V_{i}\in\mathcal{V}_{i}$. Then there exists $U_{i}\in\mathcal{U}_{i}$ such that $V_{i}$ is a subspace of $U_{i}$ and the dichotomy estimate which holds on $U_{i}$ also holds on $V_{i}$. ∎
In Remark 3 it was pointed out that for Bohl and exponential dichotomies the estimates are uniform on all one-dimensional subspaces of the corresponding spaces of the splitting. This property also holds for the new dichotomy notion defined in Definition 4.
Remark 7 (Dichotomy is uniform on one-dimensional subspaces).
If system (1) has a dichotomy on $(L_{1},L_{2})$ uniformly on the subspaces in $(\mathcal{U}_{1},\mathcal{U}_{2})$, then it also has a dichotomy on $(L_{1},L_{2})$ uniformly on the subspaces in $(\mathcal{G}_{1}(L_{1}),\mathcal{G}_{1}(L_{2}))$. This is due to the fact that if $L\neq\{0\}$ is a subspace of $\mathbb{R}^{d}$ and $\mathcal{U}\subseteq\mathcal{G}(L)$ is a cover of $L$ then
$\mathcal{G}_{1}(L)\subseteq\bigcup_{U\in\mathcal{U}}\mathcal{G}(U).$
As a consequence of Remark 5, if system (1) has a dichotomy on $(L_{1},L_{2})$ then it also has a Bohl dichotomy on $(L_{1},L_{2})$.
The notions of exponential and Bohl dichotomy spectrum (see e.g. [20]) are based on dichotomies of the $\gamma$-shifted systems for $\gamma\in\mathbb{R}$
$x(n+1)=e^{-\gamma}A(n)x(n),\qquad n\in\mathbb{T}.$ (5)
Theorem 8 (Exponential and Bohl dichotomy spectrum).
The Bohl dichotomy spectrum of (1)
[…]
as well as the exponential dichotomy spectrum of (1)
[…]
is the nonempty union of at most $d$ compact intervals.
In this paper we generalize Theorem 8 to dichotomies which hold uniformly on subspaces of prescribed dimensions and provide formulas for the endpoints of the spectral intervals (Theorems 29 and 30) which are also new for the exponential and Bohl dichotomy spectrum.
2 Dichotomies
In this section we study dichotomies which are uniform on subspaces of a splitting. The estimates (3) and (4) are the building blocks of Definition 4 which we abbreviate for convenience in the following definition.
Definition 9 (Dichotomy estimates).
Let $\gamma\in\mathbb{R}$ and $U\in\mathcal{G}(\mathbb{R}^{d})$.
  1. (a)
    We say $\mathrm{D}_{1}(\gamma,U)$ holds if there exists $C\colon U\to\mathbb{R}_{>0}$, such that
    $\|x(n,x_{0})\|\leq C(x_{0})\mathrm{e}^{\gamma(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{T},\,n\geq m,\,x_{0}\in U.$
    If the mapping $C$ is constant, we say that $\mathrm{D}_{1}(\gamma,U)$ holds uniformly.
  2. (b)
    We say $\mathrm{D}_{2}(\gamma,U)$ holds if there exists $C\colon U\to\mathbb{R}_{>0}$, such that
    $\|x(n,x_{0})\|\geq C(x_{0})\mathrm{e}^{\gamma(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{T},\,n\geq m,\,x_{0}\in U.$
    If the mapping $C$ is constant, we say that $\mathrm{D}_{2}(\gamma,U)$ holds uniformly.
Using the notation of Definition 9 system (1) has a dichotomy with splitting $(L_{1},L_{2})$, if and only if there exists $\alpha>0$ and a covering $\mathcal{U}_{1}$ of $L_{1}$ and a covering $\mathcal{U}_{2}$ of $L_{2}$, such that for all $U_{1}\in\mathcal{U}_{1}$ and $U_{2}\in\mathcal{U}_{2}$ the properties $\mathrm{D}_{1}(-\alpha,U_{1})$ and $\mathrm{D}_{2}(\alpha,U_{2})$ hold uniformly. In particular, in this case $C$ can be chosen as a constant function.
Remark 10 (Characterization of dichotomy of $\gamma$-shifted system).
For $\gamma\in\mathbb{R}$, the solution of the $\gamma$-shifted system (5) which starts at 0 in $x_{0}\in\mathbb{R}^{d}$ is $\big(\mathrm{e}^{-\gamma n}x(n,x_{0})\big)_{n\in\mathbb{T}}$. The $\gamma$-shifted system has a dichotomy with splitting $(L_{1},L_{2})$ if and only if there exists $\alpha>0$ and a covering $\mathcal{U}_{1}$ of $L_{1}$ and a covering $\mathcal{U}_{2}$ of $L_{2}$, such that $\mathrm{D}_{1}(\gamma-\alpha,U)$ holds uniformly for all $U\in\mathcal{U}_{1}$ and $\mathrm{D}_{2}(\gamma+\alpha,U)$ holds uniformly for all $U\in\mathcal{U}_{2}$.
Proposition 11 (Dichotomy estimates are dichotomous).
Let $\alpha>0$, $\gamma\in\mathbb{R}$ and $U_{1},U_{2}\in\mathcal{G}(\mathbb{R}^{d})$. If $\mathrm{D}_{1}(\gamma-\alpha,U_{1})$ and $\mathrm{D}_{2}(\gamma+\alpha,U_{2})$ hold, then $U_{1}\cap U_{2}=\{0\}$.
More generally, suppose that $U\in\mathcal{G}(\mathbb{R}^{d})$, $\alpha>0$ and $\gamma\in\mathbb{R}$. If $\mathrm{D}_{2}(\gamma+\alpha,U)$ holds and if $\big(\mathrm{e}^{-\gamma n}x(n,x_{0})\big)_{n\in\mathbb{N}}$ is bounded for all $x_{0}\in U$, then $U=\{0\}$.
Proof.
Let $x_{0}\in U_{1}\cap U_{2}$. From $\mathrm{D}_{1}(\gamma-\alpha,U_{1})$ (with $m=0$ and noting that $\mathbb{N}\subseteq\mathbb{T}$), we conclude that there exists $C_{1}(x_{0})>0$ with
$\|x(n,x_{0})\|\leq C_{1}(x_{0})\mathrm{e}^{(\gamma-\alpha)n}\|x_{0}\|,\qquad n\in\mathbb{N}.$
In particular, $\big(\mathrm{e}^{-\gamma n}x(n,x_{0})\big)_{n\in\mathbb{N}}$ is bounded. The estimate $\mathrm{D}_{2}(\gamma+\alpha,U_{2})$ yields a $C_{2}(x_{0})>0$ with
$\|x_{0}\|\leq C_{2}(x_{0})\mathrm{e}^{-(\gamma+\alpha)n}\|x(n,x_{0})\|,\qquad n\in\mathbb{N}.$
Letting $n\to\infty$, using the fact that $\alpha>0$ and the boundedness of $\big(\mathrm{e}^{-\gamma n}x(n,x_{0})\big)_{n\in\mathbb{N}}$, we conclude $x_{0}=0$. ∎
Proposition 12 (Monotonicity of dichotomy subspaces).
Let $\gamma,\widetilde{\gamma}\in\mathbb{R}$ with $\gamma\leq\widetilde{\gamma}$. Suppose that the $\gamma$-shifted system (5) has a dichotomy with splitting $(L_{1},L_{2})$ and that the $\widetilde{\gamma}-$shifted system has a dichotomy with splitting $(\widetilde{L}_{1},\widetilde{L}_{2})$. Then $L_{1}\subseteq\widetilde{L}_{1}$. If $\mathbb{T}=\mathbb{Z}$, then $L_{2}\supseteq\widetilde{L}_{2}$.
Proof.
We prove $L_{1}\subseteq\widetilde{L}_{1}$. Let $x_{1}\in L_{1}$. There are $\widetilde{x}_{1}\in\widetilde{L}_{1}$, $\widetilde{x}_{2}\in\widetilde{L}_{2}$ with $x_{1}=\widetilde{x}_{1}+\widetilde{x}_{2}$. We show $\widetilde{x}_{2}=0$ by applying Proposition 11 to the $\widetilde{\gamma}-$shifted system, so that $x_{1}=\widetilde{x}_{1}$, i.e. $L_{1}\subseteq\widetilde{L}_{1}$. Indeed, there are $\alpha_{1},\alpha_{2}>0$, such that $D_{1}(\gamma-\alpha_{1},\operatorname{span}\{x_{1}\})$ and $D_{1}(\widetilde{\gamma}-\alpha_{2},\operatorname{span}\{\widetilde{x}_{1}\})$ and $\mathrm{D}_{2}(\widetilde{\gamma}+\alpha_{2},\operatorname{span}\{\widetilde{x}_{2}\})$ hold by Remark 7. In particular, for $\alpha\coloneqq\min(\alpha_{1},\alpha_{2})$ there is $C>0$, such that for $n\in\mathbb{N}$,
$\|\mathrm{e}^{-\widetilde{\gamma}n}x(n,\widetilde{x}_{2})\|$ $=\mathrm{e}^{-\widetilde{\gamma}n}\|x(n,x_{1}-\widetilde{x}_{1})\|$
$\leq\mathrm{e}^{-\widetilde{\gamma}n}\big(C\mathrm{e}^{(\gamma-\alpha)(n-0)}\|x(0,x_{1})\|+C\mathrm{e}^{(\widetilde{\gamma}-\alpha)(n-0)}\|x(0,\widetilde{x}_{1})\|\big)$
$=C\big(\mathrm{e}^{(\gamma-\widetilde{\gamma}-\alpha)n}\|x_{1}\|+\mathrm{e}^{-\alpha n}\|\widetilde{x}_{1}\|\big).$
Since $\gamma-\widetilde{\gamma}\leq 0$, we see that $\big(\mathrm{e}^{-\widetilde{\gamma}n}x(n,\widetilde{x}_{2})\big)_{n\in\mathbb{N}}$ is bounded, so that Proposition 11 yields $\operatorname{span}\{\widetilde{x}_{2}\}=\{0\}$, i.e. $\widetilde{x}_{2}=0$. That $L_{2}\supseteq\widetilde{L}_{2}$ if $\mathbb{T}=\mathbb{Z}$, can be shown similarly. ∎
Proposition 13 (Uniqueness and dynamic characterization of dichotomy subspaces).
Let $\gamma\in\mathbb{R}$. Suppose that the $\gamma$-shifted system (5) has a dichotomy with splitting $(L_{1},L_{2})$. Then
$L_{1}=\{x_{0}\in\mathbb{R}^{d}:\lim_{n\to\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\},$
and if $\mathbb{T}=\mathbb{Z}$, then
$L_{2}=\big\{x_{0}\in\mathbb{R}^{d}:\lim_{n\to-\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\big\}.$
Moreover, if for $\widetilde{\gamma}\in\mathbb{R}$, the $\widetilde{\gamma}-$shifted system has a dichotomy with splitting $(\widetilde{L}_{1},\widetilde{L}_{2})$ and $\dim L_{1}=\dim\widetilde{L}_{1}$, then $L_{1}=\widetilde{L}_{1}$ and if $\mathbb{T}=\mathbb{Z}$, then $L_{2}=\widetilde{L}_{2}$.
Proof.
We prove the representation of $L_{1}$ only and set $M\coloneqq\{x_{0}\in\mathbb{R}^{d}:\lim_{n\to\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\}$. $L_{1}\subseteq M$, since there is $\alpha>0$, such that $D_{1}(\gamma-\alpha,\operatorname{span}\{x_{0}\})$ holds for all $x_{0}\in L_{1}$. To show $M\subseteq L_{1}$, let $x_{0}\in M$ and let $x_{1}\in L_{1}$, $x_{2}\in L_{2}$ with $x_{0}=x_{1}+x_{2}$. Similar to the proof of $\widetilde{x}_{2}=0$ in the proof of Proposition 12, it follows that $x_{2}=0$ and thus $x_{0}=x_{1}\in L_{1}$. Now suppose $\widetilde{\gamma}\in\mathbb{R}$ and assume w.l.o.g. $\widetilde{\gamma}\leq\gamma$. From Proposition 12, we obtain that $\widetilde{L}_{1}\subseteq L_{1}$ and since $\dim\widetilde{L}_{1}=\dim L_{1}$, also $\widetilde{L}_{1}=L_{1}$. Similarly one can show that $\widetilde{L}_{2}=L_{2}$ if $\mathbb{T}=\mathbb{Z}$. ∎
3 Bohl exponents and dichotomy estimates
A useful tool for studying properties of dichotomies is the concept of Bohl exponents which were introduced for individual solutions in [17] and later extended to subspace exponents in [16] and [20]. In this section we recall the upper and lower Bohl exponents from [20] and investigate their relation to the dichotomy estimates in Definition 9. We set $\sup\emptyset\coloneqq$-$\infty$ and $\inf\emptyset\coloneqq\infty$.
Definition 14 (Upper and lower Bohl exponent).
For $U\in\mathcal{G}(\mathbb{R}^{d})$ we define the upper Bohl exponent
$\overline{\beta}(U)\coloneqq\inf_{N\in\mathbb{N}}\sup\bigg\{\frac{1}{n-m}\ln\frac{\|x(n,x_{0})\|}{\|x(m,x_{0})\|}:m,n\in\mathbb{T},\,n-m>N,\,x_{0}\in U\setminus\{0\}\bigg\}$
and the lower Bohl exponent
$\underline{\beta}(U)\coloneqq\sup_{N\in\mathbb{N}}\inf\bigg\{\frac{1}{n-m}\ln\frac{\|x(n,x_{0})\|}{\|x(m,x_{0})\|}:m,n\in\mathbb{T},\,n-m>N,\,x_{0}\in U\setminus\{0\}\bigg\}.$
It should be noted that in the literature a series of other but equivalent definitions of these exponents can be found and sometimes they appear under different names (see [20, Remark 8]).
Remark 15 (Basic properties of Bohl exponents).
(a) $\overline{\beta}(\{0\})=$-$\infty$ and $\underline{\beta}(\{0\})=\infty$.
(b) If $U\neq\{0\}$ then $\ln\|A^{-1}\|_{\infty}\leq\underline{\beta}(U)\leq\overline{\beta}(U)\leq\ln\|A\|_{\infty}$. In particular,
$\underline{\beta}(U),\overline{\beta}(U)\in\mathbb{R},\qquad U\in\mathcal{G}(\mathbb{R}^{d})\setminus\{0\}.$
(c) For $U,V\in\mathcal{G}(\mathbb{R}^{d})$
$U\subseteq V\quad\Rightarrow\quad\underline{\beta}(U)\geq\underline{\beta}(V)\text{ and }\overline{\beta}(U)\leq\overline{\beta}(V).$
The next two propositions state the relation between the upper and lower Bohl exponents and the dichotomy estimates $D_{1}(\gamma,U)$ and $D_{2}(\gamma,U)$, respectively.
Proposition 16 (Upper Bohl exponent and dichotomy estimate $\mathrm{D}_{1}(\gamma,U)$).
Let $\mathcal{V}\subseteq\mathcal{G}(\mathbb{R}^{d})$ be non-empty and for each $V\in\mathcal{V}$ let $\mathcal{U}(V)\subseteq\mathcal{G}(\mathbb{R}^{d})$ be an arbitrary non-empty set of subspaces of $\mathbb{R}^{d}$. Let $\gamma\in\mathbb{R}$.
  1. (a)
    If there is $V\in\mathcal{V}$, such that for all $U\in\mathcal{U}(V)$ the estimate $\mathrm{D}_{1}(\gamma,U)$ holds uniformly, then $\gamma\geq\inf_{V\in\mathcal{V}}\sup_{U\in\mathcal{U}(V)}\overline{\beta}(U)$.
  2. (b)
    If $\gamma>\inf_{V\in\mathcal{V}}\sup_{U\in\mathcal{U}(V)}\overline{\beta}(U)$, then there is $V\in\mathcal{V}$, such that for all $U\in\mathcal{U}(V)$ the estimate $\mathrm{D}_{1}(\gamma,U)$ holds uniformly.
Proof.
We prove (a). Let $V\in\mathcal{V}$, such that $\mathrm{D}_{1}(\gamma,U)$ holds for all $U\in\mathcal{U}(V)$. First consider $U\in\mathcal{U}(V)\setminus\{0\}$. We obtain by $\mathrm{D}_{1}(\gamma,U)$ a constant $C>0$, such that
$\frac{1}{n-m}\ln\frac{\|x(n,x_{0})\|}{\|x(m,x_{0})\|}\leq\frac{1}{n-m}C+\gamma,\qquad m,n\in\mathbb{T},\,n>m,x_{0}\in U\setminus\{0\}.$
Hence for all $N\in\mathbb{N}$, we obtain
$\frac{1}{n-m}\ln\frac{\|x(n,x_{0})\|}{\|x(m,x_{0})\|}\leq\frac{C}{N}+\gamma,\qquad m,n\in\mathbb{T},\,n-m>N,x_{0}\in U\setminus\{0\}.$
Thus $\overline{\beta}(U)\leq\gamma$. Also $\overline{\beta}(\{0\})=-\infty<\gamma$ and we obtain $\sup_{U\in\mathcal{U}(V)}\overline{\beta}(U)\leq\gamma$.
We prove (b). From $\gamma>\inf_{V\in\mathcal{V}}\sup_{U\in\mathcal{U}(V)}\overline{\beta}(U)$ we conclude that there exists $V\in\mathcal{V}$, such that for all $U\in\mathcal{U}(V)$ we have
$\gamma>\overline{\beta}(U)=\inf_{N\in\mathbb{N}}\sup\bigg\{\frac{1}{n-m}\ln\frac{\|x(n,x_{0})\|}{\|x(m,x_{0})\|}:m,n\in\mathbb{T},\,n-m>N,\,x_{0}\in U\setminus\{0\}\bigg\}.$
Note that $\mathrm{D}_{1}(\gamma,\{0\})$ always holds uniformly and for $U\in\mathcal{U}(V)\setminus\{0\}$, there is $N\in\mathbb{N}$ with
$\gamma>\frac{1}{n-m}\ln\frac{\|x(n,x_{0})\|}{\|x(m,x_{0})\|},\qquad m,n\in\mathbb{T},\,n-m>N,\,x_{0}\in U\setminus\{0\}.$
Rearranging yields $\|x(n,x_{0})\|\leq\mathrm{e}^{\gamma(n-m)}\|x(m,x_{0})\|$ for $m,n\in\mathbb{T}$, nm>Nn-m>N, $x_{0}\in U$. To conclude, we have to show that there is $C>0$, such that
$\|x(n,x_{0})\|\leq C\mathrm{e}^{\gamma(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{T},\,n-m\leq N,\,x_{0}\in U.$
Indeed, for $x_{0}\in U\setminus\{0\}$ this follows from
$\mathrm{e}^{-\gamma(n-m)}\frac{\|x(n,x_{0})\|}{\|x(m,x_{0})\|}$ $=\mathrm{e}^{-\gamma(n-m)}\frac{\|A(n-1)\cdots A(m)A(m-1)\cdots A(0)x_{0}\|}{\|A(m-1)\cdots A(0)x_{0}\|}$
$\leq\mathrm{e}^{-\gamma(n-m)}\|A(n-1)\cdots A(m)\|$
and by noting that $\mathrm{e}^{-\gamma(n-m)}\|A(n-1)\cdots A(m)\|$ is bounded on $\{(m,n)\in\mathbb{T}^{2}:0\leq n-m\leq N\}$, since $A$ is bounded. Hence the inequality holds with $C:=\|A\|_{\infty}^{N}$. ∎
Proposition 17 (Lower Bohl exponent and $\mathrm{D}_{2}(\gamma,U)$).
Let $\mathcal{V}\subseteq\mathcal{G}(\mathbb{R}^{d})$ be non-empty and for every $V\in\mathcal{V}$ let $\mathcal{U}(V)\subseteq\mathcal{G}(\mathbb{R}^{d})$ be an arbitrary non-empty set of subspaces of $\mathbb{R}^{d}$. Let $\gamma\in\mathbb{R}$.
  1. (a)
    If there is $V\in\mathcal{V}$, such that for all $U\in\mathcal{U}(V)$ the estimate $\mathrm{D}_{2}(\gamma,U)$ holds uniformly, then $\gamma\leq\sup_{V\in\mathcal{V}}\inf_{U\in\mathcal{U}(V)}\underline{\beta}(U)$.
  2. (b)
    If $\gamma<\sup_{V\in\mathcal{V}}\inf_{U\in\mathcal{U}(V)}\underline{\beta}(U)$, then there is $V\in\mathcal{V}$, such that for all $U\in\mathcal{U}(V)$ the estimate $\mathrm{D}_{2}(\gamma,U)$ holds uniformly.
Proof.
The proof is along the lines of Proposition 16. ∎
4 The spectral theorem
In this section we generalize Theorem 8 and provide formulas for the spectrum. For each $\gamma\in\mathbb{R}$ which is contained in the Bohl dichotomy resolvent set of (1)
[…]
or the exponential dichotomy resolvent set
[…]
the $\gamma$-shifted system (5) admits a dichotomy on a splitting $(L_{1},L_{2})$. The subspace $L_{1}$ and, in particular, its dimension $k:=\dim L_{1}$ is unique for a fixed $\gamma$ in the resolvent set (Proposition 13).
Covers $\mathcal{U}_{1}\subseteq\mathcal{G}(L_{1})$, $\mathcal{U}_{2}\subseteq\mathcal{G}(L_{2})$ of $L_{1}$ and $L_{2}$, respectively, on which the dichotomy estimates hold uniformly (cf. Definition 4) are
$\mathcal{U}_{1}\coloneqq\mathcal{G}_{1}(L_{1}),\quad\mathcal{U}_{2}\coloneqq\mathcal{G}_{1}(L_{2})\qquad\text{if }\gamma\in\rho_{\mathrm{BD}}(A)$
and
$\mathcal{U}_{1}\coloneqq\mathcal{G}_{k}(L_{1})=\{L_{1}\},\quad\mathcal{U}_{2}\coloneqq\mathcal{G}_{d-k}(L_{2})=\{L_{2}\}\qquad\text{if }\gamma\in\rho_{\mathrm{ED}}(A).$
In both cases
$\mathcal{U}_{1}=\mathcal{G}_{j_{1}}(L_{1}),\quad\mathcal{U}_{2}=\mathcal{G}_{j_{2}}(L_{2})$ (6)
with $(j_{1},j_{2})\in\{(1,1),(k,d-k)\}$. We extend this idea and answer in the following proposition for $k\in\{0,\dots,d\}$ the question which $(j_{1},j_{2})\in\mathbb{N}\times\mathbb{N}$ have the property that for each splitting $(L_{1},L_{2})$ with $\dim L_{1}=k$ the spaces $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ in (6) are covers of $L_{1}$ and $L_{2}$, respectively (cf. Definition 4).
Definition 18 ($k$-admissible pair).
Let $k,j_{1},j_{2}\in\{0,\dots,d\}$. We say that $(j_{1},j_{2})$ is $k$-admissible if one of the following holds:
(a) If $k=0$, then $j_{1}=0$ and $j_{2}\in\{1,\dots,d\}$.
(b) If $k\in\{1,\dots,d-1\}$, then $j_{1}\in\{1,\dots,k\}$ and $j_{2}\in\{1,\dots,d-k\}$.
(c) If $k=d$, then $j_{1}\in\{1,\dots,d\}$ and $j_{2}=0$
Proposition 19 ($k$-admissible pairs and covers of splittings).
Let $k,j_{1},j_{2}\in\{0,\dots,d\}$. Then $(j_{1},j_{2})$ is $k$-admissible if and only if for each splitting $(L_{1},L_{2})$ with $\dim L_{1}=k$, the spaces $L_{1}$ and $L_{2}$ are covered by their $j_{1}$- and $j_{2}$-dimensional subspaces, respectively, i.e.
$\bigcup_{U\in\mathcal{G}_{j_{1}}(L_{1})}U=L_{1}\qquad\text{and}\qquad\bigcup_{U\in\mathcal{G}_{j_{2}}(L_{2})}U=L_{2}.$
Proof.
This follows from $\mathcal{G}_{j}(L)=\emptyset$ for $L\in\mathcal{G}(\mathbb{R}^{d})$ if $\dim L<j$. ∎
Definition 20 (Dichotomy with uniformity dimensions).
Let $(j_{1},j_{2})\in\mathbb{N}\times\mathbb{N}$. We say that system (1) has a dichotomy with uniformity dimensions $(j_{1},j_{2})$ on $(L_{1},L_{2})$ if system (1) has a dichotomy on $(L_{1},L_{2})$ uniformly on all subspaces of dimensions $(j_{1},j_{2})$, i.e. with
$\mathcal{U}_{1}=\mathcal{G}_{j_{1}}(L_{1})\qquad\text{and}\qquad\mathcal{U}_{2}=\mathcal{G}_{j_{2}}(L_{2}).$
The following remark is a consequence of the refinement Theorem 6 applied to dichotomies which are uniform on all subspaces of prescribed dimensions $(j_{1},j_{2})$.
Remark 21 (Refining uniformity dimensions).
If system (1) has a dichotomy with uniformity dimensions $(j_{1},j_{2})$ on $(L_{1},L_{2})$, then $(j_{1},j_{2})$ is $k$-admissible with $k=\dim L_{1}$ and system (1) has also a dichotomy with uniformity dimensions $(\ell_{1},\ell_{2})$ for each $k$-admissible $(\ell_{1},\ell_{2})$ which satisfies $\ell_{1}\leq j_{1}$ and $\ell_{2}\leq j_{2}$.
We are now ready to define a notion of spectrum based on dichotomies with $k$-admissible uniformity dimensions $(j_{1k},j_{2k})$ for the $\gamma$-shifted system (5) where $k$ is the dimension of the dynamically characterized set (Proposition 13)
$\{x_{0}\in\mathbb{R}^{d}:\lim_{n\to\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\}.$
The exponential and Bohl dichotomy spectrum are special cases for specific choices of uniformity dimensions (Remark 23).
Definition 22 (Dichotomy resolvent and spectrum).
For each $k\in\{0,\dots,d\}$ let $(j_{1k},j_{2k})$ be $k$-admissible. We call $J\coloneqq\big((j_{10},j_{20}),\dots,(j_{1d},j_{2d})\big)\in(\mathbb{N}\times\mathbb{N})^{d+1}$ admissible uniformity dimensions. The dichotomy resolvent of (1) with admissible uniformity dimensions $J$ is defined as
[…]
$\quad\text{uniform on subspaces with dimension $(j_{1k},j_{2k})$, where $k=\dim L_{1}$}\}$
and the dichotomy spectrum of system (1) with admissible uniformity dimensions $J$ is
$\Sigma_{J}(A)\coloneqq\mathbb{R}\setminus\rho_{J}(A).$
Remark 23 (Special cases of Bohl and exponential dichotomy spectrum).
Using Remark 3 on the uniformity subspaces of Bohl and exponential dichotomies it follows that
(a) $\Sigma_{\mathrm{BD}}(A)=\Sigma_{J}(A)$ for $J=\big((0,1),(1,1),\dots,(1,1),(1,0)\big)$.
(b) $\Sigma_{\mathrm{ED}}(A)=\Sigma_{J}(A)$ for $J=\big((0,d),(1,d-1),\dots,(d,0)\big)$.
In [20, Remark 35] it was shown that in general $\Sigma_{\mathrm{BD}}(A)\subsetneq\Sigma_{\mathrm{ED}}(A)$. For an explicit construction of an AA with this strict inclusion, see e.g. [9, 4].
Remark 24 (Resolvent set characterized by dichotomy estimates).
For admissible uniformity dimensions $J$ the following two statements are equivalent:
(i) $\gamma\in\rho_{J}(A)$,
(ii) there exists a splitting $(L_{1},L_{2})$ of $\mathbb{R}^{d}$ and $\alpha>0$, such that with $k\coloneqq\dim L_{1}$ the dichotomy estimate $D_{1}(\gamma-\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{1k}}(L_{1})$ and $D_{2}(\gamma+\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{2k}}(L_{2})$.
We now define exponents, called limiting Bohl exponents, which turn out to be the boundary points of the dichotomy spectrum. Such descriptions were previously known for the uniform exponential dichotomy only in the one-dimensional case [38] and in the multidimensional case only for the largest and smallest elements of the spectrum [20, Remark 23].
Definition 25 (Limiting Bohl exponents).
For $j,k\in\mathbb{N}$ define
$\overline{\beta}_{k,j}$ $\coloneqq\inf_{L\in\mathcal{G}_{k}(\mathbb{R}^{d})}\sup_{U\in\mathcal{G}_{j}(L)}\overline{\beta}(U),$
$\underline{\beta}_{k,j}$ $\coloneqq\sup_{L\in\mathcal{G}_{d-k}(\mathbb{R}^{d})}\inf_{U\in\mathcal{G}_{j}(L)}\underline{\beta}(U).$
Remark 26.
Let $k\in\{0,\dots,d\}$ and $(j_{1},j_{2})$ be $k$-admissible. From Remark 15 it follows that for $k=0$, resp. $k=d$,
$\overline{\beta}_{k,j_{1}}=\overline{\beta}_{0,0}=-\infty\qquad\text{resp.}\qquad\underline{\beta}_{k,j_{2}}=\underline{\beta}_{d,0}=\infty.$
From Remark 15 it also follows that
$\overline{\beta}_{k,j_{1}}\in\big[-\ln\|A^{-1}\|_{\infty},\ln\|A\|_{\infty}\big],$ $k\in\{1,\dots,d\},$
$\underline{\beta}_{k,j_{2}}\in\big[-\ln\|A^{-1}\|_{\infty},\ln\|A\|_{\infty}\big],$ $k\in\{0,\dots,d-1\}.$
Lemma 27 (Characterization of dichotomy).
Let $k\in\{0,\dots,d\}$ and $(j_{1},j_{2})$ be $k$-admissible. Then the following two statements are equivalent:
(i) $\gamma\in\big(\overline{\beta}_{k,j_{1}},\underline{\beta}_{k,j_{2}}\big)$,
(ii) there exists a splitting $(L_{1},L_{2})$ of $\mathbb{R}^{d}$ with $\dim L_{1}=k$, such that the $\gamma$-shifted system (5) has a dichotomy on $(L_{1},L_{2})$ with uniformity dimensions $(j_{1},j_{2})$.
If (i) and (ii) hold, then $L_{1}$ is independent of $\gamma\in\big(\overline{\beta}_{k,j_{1}},\underline{\beta}_{k,j_{2}}\big)$. In case $\mathbb{T}=\mathbb{Z}$, also $L_{2}$ is independent of $\gamma\in\big(\overline{\beta}_{k,j_{1}},\underline{\beta}_{k,j_{2}}\big)$.
Proof.
(i) $\Rightarrow$ (ii). Let $\gamma\in\big(\overline{\beta}_{k,j_{1}},\underline{\beta}_{k,j_{2}}\big)$. There exists $\alpha>0$ with $\gamma-\alpha>\overline{\beta}_{k,j_{1}}$ and $\gamma+\alpha<\underline{\beta}_{k,j_{2}}$. Applying Proposition 16(b) (where $\mathcal{V}$ is set to $\mathcal{G}_{k}(\mathbb{R}^{d})$ and $\mathcal{U}(V)$ is set to $\mathcal{G}_{j_{1}}(V)$ for $V\in\mathcal{V}$), we obtain the existence of a space $L_{1}\in\mathcal{G}_{k}(\mathbb{R}^{d})$, such that $D_{1}(\gamma-\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{1}}(L_{1})$. Applying Proposition 17(b) (where $\mathcal{V}$ is set to $\mathcal{G}_{d-k}(\mathbb{R}^{d})$ and $\mathcal{U}(V)$ is set to $\mathcal{G}_{j_{2}}(V)$ for $V\in\mathcal{V}$), we obtain the existence $L_{2}\in\mathcal{G}_{d-k}(\mathbb{R}^{d})$, such that $D_{2}(\gamma+\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{2}}(L_{2})$. From Proposition 11, we conclude $L_{1}\cap L_{2}=\{0\}$, so that $(L_{1},L_{2})$ is a splitting of $\mathbb{R}^{d}$ with $\dim L_{1}=k$. With Remark 24 (ii) follows.
(ii) $\Rightarrow$ (i). Suppose that $(L_{1},L_{2})$ is a splitting of $\mathbb{R}^{d}$ with $\dim L_{1}=k$, such that the $\gamma$-shifted system (5) has a dichotomy on $(L_{1},L_{2})$ with uniformity dimensions $(j_{1},j_{2})$. By Remark 24, there is $\alpha>0$, such that $D_{1}(\gamma-\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{1}}(L_{1})$ and $D_{2}(\gamma+\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{2}}(L_{2})$ for some $\alpha>0$. Propositions 16(a) and 17(a) prove that $\gamma\in\big(\overline{\beta}_{k,j_{1k}},\underline{\beta}_{k,j_{2k}}\big)$.
To show that $L_{1}$ is independent of $\gamma\in\big(\overline{\beta}_{k,j_{1}},\underline{\beta}_{k,j_{2}}\big)$, let $\gamma,\gamma^{\prime}\in\big(\overline{\beta}_{k,j_{1}},\underline{\beta}_{k,j_{2}}\big)$ with $\gamma<\gamma^{\prime}$ and let $(L_{1},L_{2})$ and $(L_{1}^{\prime},L_{2}^{\prime})$ be the respective splittings. By Proposition 13
$L_{1}=\{x_{0}\in\mathbb{R}^{d}:\lim_{n\to\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\},$
and
$L_{1}^{\prime}=\{x_{0}\in\mathbb{R}^{d}:\lim_{n\to\infty}\mathrm{e}^{-\gamma^{\prime}n}x(n,x_{0})=0\}.$
Since $\gamma<\gamma^{\prime}$, it follows that $L_{1}\subseteq L_{1}^{\prime}$. Using the fact that $\dim L_{1}=\dim L_{1}^{\prime}=k$, we conclude that $L_{1}=L_{1}^{\prime}$. Similarly it follows that $L_{2}=L_{2}^{\prime}$ in case $\mathbb{T}=\mathbb{Z}$. ∎
Lemma 28 (Dichotomy resolvent).
Let $J\coloneqq\big((j_{10},j_{20}),\dots,(j_{1d},j_{2d})\big)\in(\mathbb{N}\times\mathbb{N})^{d+1}$ be admissible uniformity dimensions (in the sense of Definition 22). Then
$\rho_{J}(A)$ $=\big(-\infty,\underline{\beta}_{0,j_{20}}\big)\cup{}$
$\phantom{i}\quad\big(\overline{\beta}_{1,j_{11}},\underline{\beta}_{1,j_{21}}\big)\cup\dots\cup\big(\overline{\beta}_{(d-1),j_{1(d-1)}},\underline{\beta}_{(d-1),j_{2(d-1)}}\big)\cup{}$
$\phantom{i}\quad\big(\overline{\beta}_{d,j_{1d}},\infty\big)$
and the union is ordered with
$\ln\|A^{-1}\|_{\infty}\leq\underline{\beta}_{k,j_{2k}}\leq\overline{\beta}_{k+1,j_{1(k+1)}}\leq\ln\|A\|_{\infty},\qquad k\in\{0,\dots,d-1\}.$
Proof.
From Lemma 27 we obtain
$\rho_{J}(A)=\bigcup_{k=0}^{d}\big(\overline{\beta}_{k,j_{1k}},\underline{\beta}_{k,j_{2k}}\big).$
The equality follows from Remark 26. We show for $k\in\{0,\dots,d-1\}$ that $\underline{\beta}_{k,j_{2k}}\leq\overline{\beta}_{k+1,j_{1(k+1)}}$, i.e.
$\sup_{L\in\mathcal{G}_{d-k}(\mathbb{R}^{d})}\inf_{U\in\mathcal{G}_{j_{2k}}(L)}\underline{\beta}(U)\leq\inf_{L\in\mathcal{G}_{k+1}(\mathbb{R}^{d})}\sup_{U\in\mathcal{G}_{j_{1(k+1)}}(L)}\overline{\beta}(U).$
[…]
$\inf_{U\in\mathcal{G}_{j_{2k}}(L)}\underline{\beta}(U)\leq\underline{\beta}(V)\leq\underline{\beta}(\operatorname{span}\{x_{0}\})\leq\overline{\beta}(\operatorname{span}\{x_{0}\})\leq\underline{\beta}(\widetilde{V})\leq\sup_{U\in\mathcal{G}_{j_{1(k+1)}}(\widetilde{L})}\underline{\beta}(U).$
We are now in a position to formulate our first main result on the dichotomy spectrum which generalizes Theorem 8 and provides a spectral flag or spectral filtration of dynamically characterized subspaces which are also related to the formulas for the spectrum given in our second main result Theorem 30 below.
Theorem 29 (Spectral theorem).
Let $J\in(\mathbb{N}\times\mathbb{N})^{d+1}$ be admissible. The dichotomy spectrum $\Sigma_{J}(A)$ is the nonempty union of at most $d$ compact intervals
$\Sigma_{J}(A)=[a_{1},b_{1}]\cup\dots\cup[a_{\ell},b_{\ell}],$
where $a_{1}\leq b_{1}<a_{2}\leq b_{2}<\dots<a_{\ell}\leq b_{\ell}$ and $\ell\in\{1,\dots,d\}$. Setting $b_{0}\coloneqq$-$\infty$ and $a_{\ell+1}\coloneqq\infty$, there exists a spectral filtration
$\{0\}=M_{0}\subsetneq M_{1}\subsetneq\dots\subsetneq M_{\ell}=\mathbb{R}^{d},$
of subspaces $M_{k}\subseteq\mathbb{R}^{d}$, $k\in\{0,\dots,\ell\}$ with
$M_{k}=\big\{x_{0}\in\mathbb{R}^{d}\,|\,\lim_{n\to\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\big\},\qquad\gamma\in(b_{k},a_{k+1}).$
For two-sided time $\mathbb{T}=\mathbb{Z}$, there exists a spectral decomposition
$\mathbb{R}^{d}=W_{1}\oplus\dots\oplus W_{\ell}$
into subspaces $W_{k}\subseteq\mathbb{R}^{d}$, with
$W_{k}=\big\{x_{0}\in\mathbb{R}^{d}\,|\,\lim_{n\to-\infty}\mathrm{e}^{-\gamma_{1}n}x(n,x_{0})=\lim_{n\to\infty}\mathrm{e}^{-\gamma_{2}n}x(n,x_{0})=0\big\},$
where $\gamma_{1}\in(b_{k-1},a_{k})$, $\gamma_{2}\in(b_{k},a_{k+1})$ and $k\in\{1,\dots,\ell\}$.
Proof.
We define $k_{0}\coloneqq 0$ and iteratively if $k_{i}$ is defined, we define $k_{i+1}\coloneqq\inf\{n\in\{k_{i}+1,\dots,d\}:(\overline{\beta}_{n,j_{1n}},\underline{\beta}_{n,j_{2n}})\neq\emptyset\}$, $i\in\mathbb{N}$. We define $\ell\in\{1,\dots,d\}$ as the maximal number, such that $k_{\ell}\neq\infty$. Then $k_{\ell}=d$ and we set
$a_{i}\coloneqq\underline{\beta}_{k_{i-1},j_{2k_{i-1}}},\qquad i\in\{1,\dots,\ell+1\}\quad\text{and}\quad b_{i}\coloneqq\overline{\beta}_{k_{i},j_{1k_{i}}},$ $i\in\{0,\dots,\ell\}.$
By Lemma 28, we obtain $a_{1}\leq b_{1}<a_{2}\leq b_{2}<\dots<a_{\ell}\leq b_{\ell}$ and
$\Sigma_{J}(A)=[a_{1},b_{1}]\cup\dots\cup[a_{\ell},b_{\ell}].$
To show the existence of the spectral filtration, choose for each $k\in\{0,\dots,\ell\}$ a $\gamma\in(b_{k},a_{k+1})$ and define
$M_{k}\coloneqq\big\{x_{0}\in\mathbb{R}^{d}\,|\,\lim_{n\to\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\big\}.$
By Proposition 13 the $\gamma$-shifted system has a dichotomy on $(L_{1},L_{2})$ with $L_{1}=M_{k}$. The fact that $M_{k}$ ist well-defined, i.e. independent of $\gamma\in(b_{k},a_{k+1})$, follows from Proposition 27. Obviously $M_{k}\subseteq M_{k+1}$ for $k\in\{0,\dots,\ell-1\}$. By Proposition 27 $\dim M_{k}<\dim M_{k+1}$, proving that the inclusion is strict.
Now suppose $\mathbb{T}=\mathbb{Z}$. Then by Proposition 13 and Lemma 27 the space
$N_{k}\coloneqq\big\{x_{0}\in\mathbb{R}^{d}:\lim_{n\to-\infty}\mathrm{e}^{-\gamma n}x(n,x_{0})=0\big\},\qquad\gamma\in(b_{k},a_{k+1}),$
is well-defined for $k\in\{0,\dots,\ell\}$, so that $W_{k}=M_{k}\cap N_{k-1}$ for $k\in\{1,\dots,\ell\}$. By Proposition 13 the $\gamma$-shifted system has a dichotomy on $(L_{1},L_{2})$ with $L_{1}=M_{k}$ and $L_{2}=N_{k}$, implying that $M_{k}\cap N_{k}=\{0\}$ and $M_{k}+N_{k}=\mathbb{R}^{d}$.
We now show the spectral decomposition. To show that the sum is direct, let $\ell,j\in\{1,\dots,d\}$ with $\ell<j$ and let $x_{0}\in(M_{\ell}\cap N_{\ell-1})\cap(M_{j}\cap N_{j-1})$. Since $M_{\ell}\subseteq M_{j-1}$, we have $x_{0}\in M_{j-1}\cap N_{j-1}=\{0\}$ by Proposition 11, proving that $x_{0}=0$. We show inductively that
$(a)$ $M_{1}=M_{1}\cap N_{0}.$
$(b)$ $\text{If $k>1$ and $M_{k-1}=\bigoplus_{j=1}^{k-1}M_{j}\cap N_{j-1}$ for $k\in\{1,\dots,d\}$},$
$\qquad\text{then $M_{k}=\bigoplus_{j=1}^{k}M_{j}\cap N_{j-1}$}.$
Part (a) follows, since $N_{0}=\mathbb{R}^{d}$. To show part (b), suppose that $k>1$ and $M_{k-1}=\bigoplus_{j=1}^{k-1}M_{j}\cap N_{j-1}$. That $\bigoplus_{j=1}^{k}M_{j}\cap N_{j-1}\subseteq M_{k}$, follows from $M_{j}\subseteq M_{k}$ for $j\in\{1,\dots,k\}$. Now suppose that $x_{0}\in M_{k}$. Let $x_{0}=x_{1}+x_{2}$, with $x_{1}\in M_{k-1}$ and $x_{2}\in N_{k-1}$. Since $x_{0}-x_{1}\in M_{k}$, we have $x_{2}=x_{0}-x_{1}\in M_{k}\cap N_{k-1}$. By assumption also $x_{1}\in M_{k-1}=\bigoplus_{j=1}^{k-1}M_{j}\cap N_{j-1}$, so that $x_{0}=x_{1}+x_{2}\in\bigoplus_{j=1}^{k}M_{j}\cap N_{j-1}$, indeed. The theorem now follows, since $M_{\ell}=\mathbb{R}^{d}$. ∎
The proof of Theorem 29 yields explicit formulas for the boundary points of the spectral intervals.
Theorem 30 (Formula for the dichotomy spectrum).
Let $J\in(\mathbb{N}\times\mathbb{N})^{d+1}$ be admissible. Let $k_{i}\coloneqq\dim M_{i}$ denote the dimension of the spectral flag subspaces $M_{i}$, $i\in\{0,\dots,\ell\}$, of Theorem 29. The following formula holds for the dichotomy spectrum $\Sigma_{J}(A)=[a_{1},b_{1}]\cup\dots\cup[a_{\ell},b_{\ell}]$ of system (1).
$a_{i}=\underline{\beta}_{k_{i-1},j_{2k_{i-1}}}\qquad\text{and}\qquad b_{i}=\overline{\beta}_{k_{i},j_{1k_{i}}},\qquad i\in\{1,\dots,\ell\}.$
Remark 31 (Formulas for the Bohl and exponential dichotomy spectrum).
Using the notation of Theorem 30, we obtain the following formulas for the boundary points of the Bohl dichotomy spectrum and the exponential dichotomy spectrum (cf. Remark 23).
(a) The boundary points of the Bohl dichotomy spectrum are
$a_{i}=\underline{\beta}_{k_{i-1},1}\qquad\text{and}\qquad b_{i}=\overline{\beta}_{k_{i},1},\qquad i\in\{1,\dots,\ell\}.$
(b) The boundary points of the exponential dichotomy spectrum are
$a_{i}=\underline{\beta}_{k_{i-1},d-k_{i-1}}\qquad\text{and}\qquad b_{i}=\overline{\beta}_{k_{i},d-k_{i}},\qquad i\in\{1,\dots,\ell\}.$
In case system (1) has a dichotomy then the limiting Bohl exponents of Definition 25 can also be expressed in terms of the splitting of the dichotomy.
Corollary 32 (Limiting Bohl exponents in case of dichotomy).
If system (1) has a dichotomy on $(L_{1},L_{2})$ with uniformity dimensions $(j_{1},j_{2})$ then
$\overline{\beta}_{k,j_{1}}=\sup_{U\in\mathcal{G}_{j_{1}}(L_{1})}\overline{\beta}(U)$ (7)
where $k=\dim L_{1}$, and if $\mathbb{T}=\mathbb{Z}$, then
$\underline{\beta}_{k,j_{2}}=\inf_{U\in\mathcal{G}_{j_{2}}(L_{2})}\overline{\beta}(U).$ (8)
Proof.
The first part follows readily from Lemma 27 and we only proof formula (7). Suppose that system (1) has a dichotomy with splitting $(L_{1},L_{2})$ uniform on subspaces of dimensions $j_{1}$ and $j_{2}$. Let $\alpha>0$ with $\alpha\in(\overline{\beta}_{k,j_{1}},0)$, so that $\mathrm{D}_{1}(-\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{1}}(L_{1})$ by Proposition 16(b). Let $k\coloneqq\dim L_{1}$. If $k=0$, then $j_{1}=0$ and (7) is satisfied (cf. Remark 15). Let $k>0$. Suppose that $L_{1}^{\prime}\in\mathcal{G}_{k}(\mathbb{R}^{d})$ with
$\sup_{U^{\prime}\in\mathcal{G}_{j_{1}}(L_{1}^{\prime})}\overline{\beta}(U^{\prime})\leq\sup_{U\in\mathcal{G}_{j_{1}}(L_{1})}\overline{\beta}(U).$ (9)
It holds that $\overline{\beta}(U)\leq-\alpha$ for all $U\in\mathcal{G}_{j_{1}}(L_{1})$ (cf. Proposition 16(a) with $\mathcal{V}\coloneqq\{L_{1}\}$ and $\mathcal{U}(L_{1})\coloneqq\mathcal{G}_{j_{1}}(L_{1})$). By the inequality (9), also $\overline{\beta}(U^{\prime})\leq-\alpha$ for all $U^{\prime}\in\mathcal{G}_{j_{1}}(L_{1}^{\prime})$. By Proposition 16(b), we conclude that $\mathrm{D}_{1}(-\alpha/2,U^{\prime})$ holds for all $U^{\prime}\in\mathcal{G}_{j_{1}}(L_{1}^{\prime})$, so that
$\underset{n\rightarrow\infty}{\lim}x(n,x_{0})=0,\qquad x_{0}\in L_{1}^{\prime}.$
This implies that $L_{1}^{\prime}\subseteq L_{1}$ by Proposition 13. But since $\dim L_{1}=\dim L_{1}^{\prime}$, also $L_{1}=L_{1}^{\prime}$. We conclude that for all $L_{1}^{\prime}\in\mathcal{G}_{k}(\mathbb{R}^{d})$, we have
$\sup_{U^{\prime}\in\mathcal{G}_{j_{1}}(L_{1}^{\prime})}\overline{\beta}(U^{\prime})\geq\sup_{U\in\mathcal{G}_{j_{1}}(L_{1})}\overline{\beta}(U).$
Taking the infimum over all $L_{1}^{\prime}\in\mathcal{G}_{k}(\mathbb{R}^{d})$ yields formula (7). ∎
Remark 33 (Open problem).
It is not clear to the authors, if (8) holds for all $L_{2}$, complementary to $L_{1}$ in the case $\mathbb{T}=\mathbb{N}$.
5 Maximal subspaces of uniformity
In this section we introduce maximality of the uniformity dimensions of a dichotomous system in Definition 34 and discuss its dependence on the dichotomy splitting in Theorems 37 and 38. An open problem and a conjecture are formulated in Remark 39.
Definition 34 (Maximal uniformity dimensions).
Suppose system (1) has a dichotomy on $(L_{1},L_{2})$. Then $(u_{1},u_{2})$ are called maximal uniformity dimensions if
(a) system (1) has a dichotomy on $(L_{1},L_{2})$ with uniformity dimensions $(u_{1},u_{2})$, and
(b) if (1) has a dichotomy on $(L_{1},L_{2})$ with uniformity dimensions $(\ell_{1},\ell_{2})$ then
$\ell_{1}\in\{0,\dots,u_{1}\}\qquad\text{and}\qquad\ell_{2}\in\{0,\dots,u_{2}\}.$
Remark 35 (Dichotomy implies existence of maximal uniformity dimensions).
If system (1) has a dichotomy on $(L_{1},L_{2})$ there exist maximal uniformity dimensions $(u_{1},u_{2})\in\{0,\dots,\dim L_{1}\}\times\{0,\dots,\dim L_{2}\}$.
In case system (1) has a dichotomy on $(L_{1},L_{2})$ then by Proposition 13 the space $L_{1}$ is unique and hence the maximal uniformity dimension $u_{1}$ does not depend on $L_{1}$. If $\mathbb{T}=\mathbb{Z}$, the space $L_{2}$ is unique and in this case the maximal uniformity dimension $u_{2}$ does not depend on $L_{2}$. We provide a partial answer to the question how $u_{2}$ does depend on $L_{2}$ in Theorem 38 and formulate an open problem in Remark 39. As preparatory results we prove the following Lemma and Theorem 37.
Lemma 36 (Dichotomy estimate follows from uniform estimate at later time).
Let $\gamma\in\mathbb{R}$, $U\in\mathcal{G}(\mathbb{R}^{d})$, $C>0$ and $m_{0}\in\mathbb{N}$ with
$\|x(n,x_{0})\|\geq C\mathrm{e}^{\gamma(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{N},\,n\geq m\geq m_{0},\,x_{0}\in U.$ (10)
Then $\mathrm{D}_{2}(\gamma,U)$ holds uniformly.
Proof.
If $m_{0}=0$, then the statement of Lemma 36 is clear. Suppose that $m_{0}\geq 1$. By induction, it suffices to show that
$\|x(n,x_{0})\|\geq C^{\prime}\mathrm{e}^{\gamma(n-m_{0}+1)}\|x(m_{0}-1,x_{0})\|,\qquad n\in\mathbb{N},\,n\geq m_{0}-1,\,x_{0}\in U$ (11)
holds for some $C^{\prime}>0$. Let us denote
$a=\max\big\{\|A\|_{\infty},\|A^{-1}\|_{\infty}\big\}.$
Considering (10) for $n$ set to $n+1$ and for $m$ set to $m_{0}$, we get
$\|x(n+1,x_{0})\|\geq C\mathrm{e}^{\gamma(n-m_{0}+1)}\|x(m_{0},x_{0})\|,\qquad m,n\in\mathbb{N},\,n\geq m_{0}-1,\,x_{0}\in U$
and consequently
$\|A^{-1}(m_{0}-1)\|\cdot\|x(n+1,x_{0})\|\geq C\mathrm{e}^{\gamma(n-m_{0}+1)}\|A^{-1}(m_{0}-1)\|\cdot\|x(m_{0},x_{0})\|.$
Noting that
$\|A^{-1}(m_{0}-1)\|\|x(n+1,x_{0})\|$ $\leq a\|x(n+1,x_{0})\|$
$=a\|A(n)x(n,x_{0})\|$
$\leq a^{2}\|x(n,x_{0})\|$
and
$\|A^{-1}(m_{0}-1)\|\|x(m_{0},x_{0})\|\geq\|A^{-1}(m_{0}-1)x(m_{0},x_{0})\|=\|x(m_{0}-1,x_{0})\|,$
we obtain
$a^{2}\|x(n,x_{0})\|\geq C\mathrm{e}^{\gamma(n-m_{0}+1)}\|x(m_{0}-1,x_{0})\|.$
This proves inequality (10) with $C^{\prime}=C/a^{2}$. ∎
Theorem 37 (Independence of maximal uniformity dimensions of splitting).
Suppose that $T=\mathbb{N}$. Let $(L_{1},L_{2})$ and $(L_{1},L_{2}^{\prime})$ be splittings of $\mathbb{R}^{d}$. Suppose that System (1) has a dichotomy on $(L_{1},L_{2})$ with maximal uniformity dimensions $(u_{1},u_{2})$ and on $(L_{1},L_{2}^{\prime})$ and with maximal uniformity dimensions $(u_{1}^{\prime},u_{2}^{\prime})$. Then
$u_{1}=u_{1}^{\prime}.$ (12)
If additionally for each $V^{\prime}\in\mathcal{G}_{u_{2}}(L_{2}^{\prime})$, we have $\dim V_{L_{1}}^{\prime}\leq u_{1}$, then
$u_{2}\leq u_{2}^{\prime}.$ (13)
Proof.
The equality (12) is clear. Let $\alpha>0$ be such that $\mathrm{D}_{1}(-\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{1}}(L_{1})$ and $\mathrm{D}_{2}(\alpha,U)$ holds uniformly for all $U\in\mathcal{G}_{j_{2}}(L_{2})$ for all $j_{1}\in\{0,...,u_{1}\}$ and $j_{2}\in\{0,...,u_{2}\}$. Let $V^{\prime}\in\mathcal{G}_{u_{2}}(L_{2}^{\prime})$. To prove the inequality (13), it suffices to show that the estimate $\mathrm{D}_{2}(\alpha,V^{\prime})$ holds uniformly. Since $\dim\pi_{L_{1}}[V^{\prime}]\leq u_{1}$ and $\dim\pi_{L_{2}}[V^{\prime}]\leq\dim V^{\prime}=u_{2}$, there exist constants $C_{1},C_{2}>0$ such that
$\|x(n,x_{0})\|\leq C_{1}\mathrm{e}^{-\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{N},\,n\geq m,\,x_{0}\in\pi_{L_{1}}[V^{\prime}]$ (14)
and
$\|x(n,x_{0})\|\geq C_{2}\mathrm{e}^{\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{N},\,n\geq m,\,x_{0}\in\pi_{L_{2}}[V^{\prime}].$ (15)
Let $K>0$ be such that
$C_{1}\leq K\qquad\text{and}\qquad C_{2}\geq K^{-1}.$
Then (14) and (15) imply that
$\|x(n,x_{0})\|\leq K\mathrm{e}^{-\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{N},\,n\geq m,\,x_{0}\in\pi_{L_{1}}[V^{\prime}]$ (16)
and
$\|x(n,x_{0})\|\geq K^{-1}\mathrm{e}^{\alpha(n-m)}\|x(m,x_{0})\|,\qquad m,n\in\mathbb{N},\,n\geq m,\,x_{0}\in\pi_{L_{2}}[V^{\prime}].$ (17)
Consider $v^{\prime}\in V^{\prime}$ with $\|v^{\prime}\|=1$ and note that $v^{\prime}\neq 0$ implies $\pi_{L_{2}}v^{\prime}\neq 0$. We have
$\|x(n,v^{\prime})\|\geq\|x(n,\pi_{L_{2}}v^{\prime})\|-\|x(n,\pi_{L_{1}}v^{\prime})\|,\qquad n\in\mathbb{N}$
and by applying (16) and (17), we get
$\|x(n,v^{\prime})\|$ $\geq K^{-1}\mathrm{e}^{\alpha(n-m)}\|x(m,\pi_{L_{2}}v^{\prime})\|-K\mathrm{e}^{-\alpha(n-m)}\|x(m,\pi_{L_{1}}v^{\prime})\|$
$\geq K^{-1}\mathrm{e}^{\alpha(n-m)}\|x(m,\pi_{L_{2}}v^{\prime})\|-K\mathrm{e}^{\alpha(n-m)}\|x(m,\pi_{L_{1}}v^{\prime})\|$
$=\mathrm{e}^{\alpha(n-m)}(K^{-1}\|x(m,\pi_{L_{2}}v^{\prime})\|-K\|x(m,\pi_{L_{1}}v^{\prime})\|),$
$\mathrm{e}^{\alpha(n-m)}(K^{-1}-K\frac{\|x(m,\pi_{L_{1}}v^{\prime})\|}{\|x(m,\pi_{L_{2}}v^{\prime})\|})\|x(m,\pi_{L_{2}}v^{\prime})\|,\qquad m,n\in\mathbb{N},\,n\geq m.$
Taking in (14) $x_{0}=\pi_{L_{1}}v^{\prime}$ and $m=0$, and in (15) $x_{0}=\pi_{L_{2}}v^{\prime}$ and $m=0$, we get
$\frac{\|x(k,\pi_{L_{1}}v^{\prime})\|}{\|x(k,\pi_{L_{2}}v^{\prime})\|}\leq\frac{K^{2}}{\mathrm{e}^{2\alpha k}}\frac{\|\pi_{L_{1}}v^{\prime}\|}{\|\pi_{L_{2}}v^{\prime}\|},\qquad k\in\mathbb{N}$ (18)
and therefore
$\|x(n,v^{\prime})\|\geq\mathrm{e}^{\alpha(n-m)}\bigg(K^{-1}-\frac{K^{3}}{\mathrm{e}^{2\alpha m}}\frac{\|\pi_{L_{1}}v^{\prime}\|}{\|\pi_{L_{2}}v^{\prime}\|}\bigg)\|x(m,\pi_{L_{2}}v^{\prime})\|,\qquad m,n\in\mathbb{N},\,n\geq m.$ (19)
The function $v^{\prime}\mapsto\frac{\|\pi_{L_{1}}v^{\prime}\|}{\|\pi_{L_{2}}v^{\prime}\|}$ defined on the compact set $\{v^{\prime}\in V:\|v^{\prime}\|=1\}$ is continuous and therefore attains its maximum value. Hence
$C^{\prime}\coloneqq\max\bigg\{\frac{\|\pi_{L_{1}}v^{\prime}\|}{\|\pi_{L_{2}}v^{\prime}\|}:v^{\prime}\in V,\,\|v^{\prime}\|=1\bigg\}\in[0,\infty),$
is well-defined. The inequality (19) then implies that
$\|x(n,v^{\prime})\|\geq\mathrm{e}^{\alpha(n-m)}(K^{-1}-\frac{K^{3}}{\mathrm{e}^{2\alpha m}}C^{\prime})\|x(m,\pi_{L_{2}}v^{\prime})\|,\qquad m,n\in\mathbb{N},\,n\geq m.$
Consider $m_{0}\in\mathbb{N}$ such that
$K^{-1}-\frac{K^{3}}{\mathrm{e}^{2\alpha m_{0}}}C^{\prime}>0.$
Define
$C^{\prime\prime}\coloneqq K^{-1}-\frac{K^{3}}{\mathrm{e}^{2\alpha m_{0}}}C^{\prime}.$
We get
$\|x(n,v^{\prime})\|\geq\mathrm{e}^{\alpha(n-m)}C^{\prime\prime}\|x(m,\pi_{L_{2}}v^{\prime})\|,\qquad m,n\in\mathbb{N},\,n\geq m\geq m_{0},$ (20)
since
$K^{-1}-\frac{K^{3}}{\mathrm{e}^{2\alpha m}}C^{\prime}\geq C^{\prime\prime}$
for $m\geq m_{0}$. Note that by (18), we have
$\|x(k,v^{\prime})\|$ $\leq\|x(k,\pi_{L_{2}}v^{\prime})\|+\|x(k,\pi_{L_{1}}v^{\prime})\|$
$\leq\|x(k,\pi_{L_{2}}v^{\prime})\|+\frac{K^{2}}{\mathrm{e}^{2\alpha k}}C^{\prime}\|x(k,\pi_{L_{2}}v^{\prime})\|$
$=(1+\frac{K^{2}}{\mathrm{e}^{2\alpha k}}C^{\prime})\|x(k,\pi_{L_{2}}v^{\prime})\|$
$\leq(1+K^{2}C^{\prime})\|x(k,\pi_{L_{2}}v^{\prime})\|,\qquad k\in\mathbb{N}$
and therefore
$\|x(m,y(v^{\prime}))\|\geq(1+K^{2}C^{\prime})^{-1}\|x(m,v^{\prime})\|,\qquad m\in\mathbb{N}.$ (21)
From the inequalities (20) and (21), we get
$\|x(n,v^{\prime})\|\geq\mathrm{e}^{\alpha(n-m)}C^{\prime\prime}(1+K^{2}C^{\prime})^{-1}\|x(m,v^{\prime})\|,\qquad m,n\in\mathbb{N},\,n\geq m\geq m_{0}.$
The statement now follows from Lemma 36. ∎
Theorem 38 (Maximal uniformity dimensions for one-sided time).
Suppose that $T=\mathbb{N}$. Let $(L_{1},L_{2})$ be a splitting of $\mathbb{R}^{d}$. Suppose that System (1) has a dichotomy on $(L_{1},L_{2})$ with maximal uniformity dimensions $(u_{1},u_{2})$. Denote by $(u_{1},u_{2}^{\perp})$ the maximal uniformity dimensions of system (1) with respect to the splitting $(L_{1},L_{1}^{\perp})$. Then it holds that $u_{2}\leq u_{2}^{\perp}$.
Proof.
This follows from Theorem 37, noting that for each $V^{\prime}\in\mathcal{G}_{u_{2}}(L_{1}^{\perp})$, we have $\dim V_{L_{1}}[V^{\prime}]=0\leq u_{1}$
In the following remark we list those cases in which Theorem 37 yields a unique maximal uniformity dimension of a dichotomy, i.e. independent of the splitting, and we formulate a conjecture on the general dependence of the maximal uniformity dimensions on the subspace $L_{2}$ which is complementary to $L_{1}$.
Remark 39 (Dependence of maximal uniformity dimensions on splitting for one-sided time).
Suppose that system (1) with $\mathbb{T}=\mathbb{N}$ has a dichotomy on a splitting $(L_{1},L_{2})$.
(a) The maximal uniformity dimensions do not depend on $L_{2}$ in case $\dim L_{1}=1$ or $\dim L_{1}=d-1$.
(b) Conjecture: There exists $A\colon\mathbb{N}\to\mathrm{GL}(\mathbb{R}^{d})$ such that the maximal uniformity dimension $u_{2}$ depends on the choice of $L_{2}$.
Acknowledgements
The research of Adam Czornik was supported by the Polish National Agency for Academic Exchange according to the decision No. BPN/BEK/2025/1/00054/DEC/1.